independent random sample of size 9 is taken with replacement from a population with mean 25.2 and variance 12


1
Expert's answer
2021-06-18T09:29:53-0400
"\mu_{\bar{x}}=\mu=25.2"

"\sigma^2_{\bar{x}}=\dfrac{\sigma^2}{n}=\dfrac{12}{9}=\dfrac{4}{3}"

"\sigma_{\bar{x}}=\dfrac{\sigma}{\sqrt{n}}=\dfrac{\sqrt{12}}{\sqrt{9}}=\dfrac{2\sqrt{3}}{3}"